Find the area of the largest isosceles triangle inscribed in a semicircle of radius 10ft., the vertex of the triangle being at the center of the circle.
1
Expert's answer
2011-08-09T12:47:12-0400
The area of isosceles triangle can be found by the following expression:
S = 1/2 a2 sin θ
The largest value of S would be then sin θ = 1, thus θ = 90 degree.
Thus
S = 1/2 102 = 50 ft2.